### v1

@@ Abstract.
* Co-first authors
copyright: none
@@ 4.1. RQ1: Overall Performance
Holistic WizardCoder-15B-V1.0 90.00 62.86 94.29 71.43 72.00 54.29 71.43 34.29 55.00 0.00 31.76 2.86 70.81 52.59 70.19 33.36 32.79 0.00 31.40 2.57
Holistic deepseek-coder-6.7b-instruct 88.00 87.14 100.00 100.00 64.00 68.57 82.86 71.43 0.00 0.00 70.59 54.29 61.59 67.87 82.55 62.21 0.00 0.00 70.59 10.55
Holistic deepseek-coder-33b-instruct 100.00 74.29 100.00 100.00 68.00 65.71 77.14 85.71 60.00 7.27 29.41 71.43 65.86 64.81 76.78 82.04 54.79 3.36 29.41 34.34
Holistic Phind-CodeLlama-34B-v2 88.00 80.00 90.00 80.00 82.00 60.00 80.00 57.14 45.00 0.00 62.35 20.00 79.34 58.76 79.66 54.58 31.03 0.00 62.32 8.42
Holistic gpt-3.5-turbo-1106 90.00 84.29 100.00 94.29 82.00 77.14 84.29 74.29 50.00 0.00 70.59 17.14 81.18 75.98 83.90 72.24 41.65 0.00 70.59 5.40
Average (Holistic) 88.73 70.61 32.38 68.81 24.26
Independent WizardCoder-15B-V1.0 76.00 67.14 70.00 25.71 52.00 60.00 47.14 14.29 0.00 0.00 0.00 0.00 51.31 58.41 46.92 13.83 0.00 0.00 0.00 0.00
Independent deepseek-coder-6.7b-instruct 66.00 90.00 97.14 85.71 64.00 67.14 78.57 62.86 0.00 0.00 70.59 42.86 61.59 66.40 78.57 55.33 0.00 0.00 70.59 12.75
Independent deepseek-coder-33b-instruct 72.00 62.86 92.86 100.00 62.00 60.00 74.29 74.29 37.50 0.00 37.65 60.00 59.61 58.57 74.26 69.45 30.37 0.00 37.54 21.07
Independent Phind-CodeLlama-34B-v2 38.00 77.14 68.57 71.43 28.00 61.43 61.43 48.57 0.00 0.00 25.88 11.43 25.73 60.56 61.30 46.96 0.00 0.00 25.88 2.71
Independent gpt-3.5-turbo-1106 90.00 88.57 91.43 71.43 76.00 81.43 72.86 42.86 12.50 0.00 64.71 0.00 74.61 80.80 72.86 42.64 12.50 0.00 64.71 0.00
Average (Independent) 75.1 59.46 18.16 57.99 13.91
Incremental WizardCoder-15B-V1.0 68.00 64.29 47.14 22.86 46.00 47.14 35.71 11.43 0.00 0.00 0.00 0.00 45.31 46.51 35.68 11.24 0.00 0.00 0.00 0.00
Incremental deepseek-coder-6.7b-instruct 70.00 85.71 98.57 91.43 70.00 68.57 82.86 68.57 0.00 0.00 70.59 34.29 67.59 67.43 82.15 64.30 0.00 0.00 70.59 8.12
Incremental deepseek-coder-33b-instruct 58.00 64.29 91.43 85.71 44.00 62.86 70.00 54.29 0.00 0.00 14.12 0.00 41.86 62.83 69.66 49.81 0.00 0.00 14.02 0.00
Incremental Phind-CodeLlama-34B-v2 44.00 75.71 65.71 54.29 38.00 61.43 58.57 42.86 0.00 0.00 22.35 5.71 35.84 60.54 57.74 41.01 0.00 0.00 21.87 2.53
Incremental gpt-3.5-turbo-1106 70.00 85.71 88.57 91.43 60.00 80.00 71.43 68.57 12.50 0.00 0.00 25.71 57.80 79.80 71.43 64.84 12.50 0.00 0.00 8.10
Average (Incremental) 71.14 57.11 9.26 55.67 6.89
Average 78.32 62.40 19.93 60.82 15.08
@@
Generally, Completion@1 of all LLMs on JavaBench was 78%, with variances when different synthesis strategies were applied. Considering a finer granularity (class-wise), the average Compilation@1 and Pass@1 were around 60% (62.40% and 60.82%, respectively). The best scores were achieved using the holistic strategy with 88.73% Completion@1, 70.61% Compilation@1, and 68.81% Pass@1. Among all LLMs, DeepSeek-Coder-33b performed the best, followed by gpt-3.5-turbo and DeepSeek-Coder-6.7b.
@@
[style=MyFrame] Finding 1: The best average of 88.73%, 70.61%, and 68.81% Completion@1, Compilation@1, and Pass@1, respectively, could be achieved on JavaBench over the studied LLMs. The Top-3 performing LLMs were DeepSeek-Coder-33b, gpt-3.5-turbo, and DeepSeek-Coder-6.7b among the studied LLMs.
Comparing the synthesis strategies, holistic synthesis was generally better than independent and incremental among all LLMs, and the declines could be significant in some cases. For example, WizardCoder dropped 48.57% (71.43% - 22.86%) on Completion@1, 31.76% at Compilation@1 and 31.40% at Completion@1 on P3. Other LLMs also observed similar drops when switching from holistic to independent or incremental synthesis. Although there were occasional cases where incremental or independent synthesis brought improvements, those were sporadic events, and the improvement was subtle, e.g., Completion@1 of WizardCoder improved 1.43% (64.29% - 62.86%) on P2. We also noticed that this observation was different from the observation of the previous work (Du et al., 2023) where they observed open-sourced LLMs performed better using independent synthesis than holistic synthesis. This could be because we used different programming languages (Python VS. Java) and code granularities (Class-level VS. Project-level).
@@
Necessity of Finer-grained Evaluation Granularity. LLMs performed similarly on four projects (P1-P4), with P1 and P3 slightly better than P2 and P4. In particular, the class-wise scores were always better than test-wise scores, with a gap up to 48.78% (= 55.67%-6.89%). This was in line with our claim in Section 3.2.1: finer-grained granularity can capture more nuanced successful cases. We also calculated the Pass@1 under project-wise granularity and found that none of the projects can be correctly completed, yielding all-0 results under all settings.
@@
In addition, we increased the size of $k$ from 1 to 5 to investigate the improvements brought by more trials. The detailed experiment results are omitted due to space limits and can be found on our website (Section 8). Overall, the average Completion/Compilation/Pass@5 under holistic strategy across all LLMs are 94.21%(+5.48%), 81.43%(+10.82%), 81.43%(+12.62%) at the class-wise granularity; 94.21%(+5.48%), 42.79%(+10.41%), 41.17%(+16.91%) at the test-wise granularity. The Pass@5 in project-wise granularity is still all-0s under all settings.
{mdframed}
[style=MyFrame] Finding 4: Increasing $k$ yields further improvement. The best average test-wise Pass@5 in JavaBench is 41.17%. Compared with 90.93% achieved by undergraduate students in project-wise evaluation (Section 2.2.3), LLMs’ capability in Java project-level programming still has much room to improve.
@@ 4.2. RQ2: Impact of Context Selection
Selected WizardCoder-15B-V1.0 90.00 62.86 94.29 71.43 72.00 54.29 71.43 34.29 55.00 0.00 31.76 2.86 70.81 52.59 70.19 33.36 32.79 0.00 31.40 2.57
Selected deepseek-coder-6.7b-instruct 88.00 87.14 100.00 100.00 64.00 68.57 82.86 71.43 0.00 0.00 70.59 54.29 61.59 67.87 82.55 62.21 0.00 0.00 70.59 10.55
Selected deepseek-coder-33b-instruct 100.00 74.29 100.00 100.00 68.00 65.71 77.14 85.71 60.00 7.27 29.41 71.43 65.86 64.81 76.78 82.04 54.79 3.36 29.41 34.34
Selected Phind-CodeLlama-34B-v2 88.00 80.00 90.00 80.00 82.00 60.00 80.00 57.14 45.00 0.00 62.35 20.00 79.34 58.76 79.66 54.58 31.03 0.00 62.32 8.42
Selected gpt-3.5-turbo-1106 90.00 84.29 100.00 94.29 82.00 77.14 84.29 74.29 50.00 0.00 70.59 17.14 81.18 75.98 83.90 72.24 41.65 0.00 70.59 5.40
Average (Selected) 88.73 70.61 32.38 68.81 24.46
Maximum WizardCoder-15B-V1.0 92.00 60.00 87.14 48.57 64.00 55.71 58.57 14.29 37.50 0.00 21.18 0.00 63.68 54.86 57.85 13.29 33.66 0.00 21.18 0.00
Maximum deepseek-coder-6.7b-instruct 98.00 91.43 100.00 88.57 74.00 70.00 71.43 80.00 37.50 0.00 29.41 68.57 73.65 68.38 71.24 68.06 32.31 0.00 29.41 13.39
Maximum deepseek-coder-33b-instruct 100.00 91.43 100.00 97.14 74.00 77.14 75.71 65.71 67.50 0.00 45.88 22.86 74.00 76.42 75.63 60.07 65.00 0.00 45.67 10.63
Maximum Phind-CodeLlama-34B-v2 82.00 88.57 97.14 88.57 68.00 60.00 74.29 54.29 12.50 0.00 58.82 20.00 67.96 59.08 74.09 50.62 12.15 0.00 58.82 7.69
Maximum gpt-3.5-turbo-1106 96.00 85.71 100.00 91.43 68.00 68.57 64.29 62.86 55.00 0.00 36.47 22.86 66.98 66.07 63.88 59.77 51.47 0.00 36.47 7.78
Average (Maximum) 89.19 65.04 26.80 63.28 21.28
Minimum WizardCoder-15B-V1.0 96.00 65.71 90.00 65.71 16.00 21.43 48.57 0.00 0.00 0.00 5.88 0.00 15.94 21.10 48.52 0.00 0.00 0.00 5.88 0.00
Minimum deepseek-coder-6.7b-instruct 86.00 88.57 100.00 85.71 52.00 15.71 57.14 22.86 5.00 0.00 0.00 0.00 52.00 15.69 57.14 19.14 5.00 0.00 0.00 0.00
Minimum deepseek-coder-33b-instruct 96.00 81.43 98.57 77.14 44.00 17.14 65.71 11.43 7.50 0.00 37.65 0.00 43.75 16.98 64.98 5.21 7.50 0.00 37.43 0.00
Minimum Phind-CodeLlama-34B-v2 82.00 87.14 97.14 82.86 46.00 44.29 57.14 8.57 0.00 0.00 0.00 0.00 46.00 43.40 57.14 8.02 0.00 0.00 0.00 0.00
Minimum gpt-3.5-turbo-1106 96.00 85.71 95.71 65.71 60.00 40.00 62.86 17.14 12.50 0.00 0.00 0.00 60.00 39.85 62.86 16.89 12.50 0.00 0.00 0.00
Average (Minimum) 86.16 35.40 3.43 34.73 3.42
Overall Average 88.02 57.02 20.87 55.61 16.39
@@
From Table 6, it is clear that among three context settings, the selected context yield the best overall results, with 68.61% (class-wise) and 24.46% (test-wise) Pass@1, which echos the results in Table 5. Though maximum context brings a 0.46% (= 89.19%-88.73%) improvement to the selected context in completion@1, the performance of using maximum context on the subsequent metrics (i.e., Compilation@1 and Pass@1) is not as good as the selected context.
@@
[style=MyFrame] Finding 5: The selected context is the best setting on JavaBench, resulting in 68.61% (class-wise) Pass@1.
@@
By combining Table 6 and Figure 4, we can make two observations. First, more context is not always a benefit. For example, a 32.31% (= 32.31%-0.00%) increase in test-wise Pass@1 is achieved by DeepSeek-6.7b when switching to the maximum context in P1, while a 23.71% (= 34.34%-10.63%) drop of test-wise Pass@1 in P4 can be observed in DeepSeek-33b. Second, only providing the class to be completed is insufficient without dependencies. We can see from Table 6 that the Pass@1 in the test-wise granularity is almost all zeros under the minimum context setting, meaning that it is almost impossible to generate project-level code that can pass test cases. In addition, the selected context includes only the method signatures (as explained in Section 3.1.1), which turned out to be more effective than the maximum context setting.

### v2

@@ Abstract.
* Co-first authors.
@@ 4.1. RQ1: Overall Performance
Holistic WizardCoder-15B-V1.0 98.00 70.00 94.29 80.00 64.00 58.57 71.43 37.14 17.50 0.00 31.76 5.71 63.99 56.87 70.19 36.14 16.51 0.00 31.40 5.03
Holistic deepseek-coder-6.7b-instruct 82.00 90.00 100.00 100.00 64.00 71.43 82.86 71.43 15.00 0.00 70.59 54.29 62.67 70.83 82.55 62.21 14.29 0.00 70.59 10.55
Holistic deepseek-coder-33b-instruct 100.00 95.71 100.00 100.00 68.00 81.43 77.14 85.71 60.00 18.18 29.41 71.43 68.00 80.49 76.78 82.04 60.00 8.74 29.41 34.34
Holistic Phind-CodeLlama-34B-v2 96.00 88.57 91.43 80.00 86.00 68.57 80.00 57.14 80.00 0.00 62.35 20.00 85.81 67.54 79.66 54.58 58.06 0.00 62.32 8.42
Holistic gpt-3.5-turbo-1106 90.00 84.29 100.00 94.29 86.00 77.14 84.29 74.29 75.00 0.00 70.59 17.14 85.87 75.98 83.90 72.24 62.29 0.00 70.59 5.40
Average (Holistic) 91.73 72.33 34.95 70.92 27.40
Independent WizardCoder-15B-V1.0 74.00 72.86 71.43 31.43 44.00 62.86 47.14 17.14 0.00 0.00 0.00 0.00 44.00 61.17 46.92 16.62 0.00 0.00 0.00 0.00
Independent deepseek-coder-6.7b-instruct 80.00 90.00 97.14 85.71 70.00 67.14 78.57 62.86 35.00 0.00 70.59 42.86 69.29 66.40 78.57 55.33 31.55 0.00 70.59 12.75
Independent deepseek-coder-33b-instruct 76.00 78.57 92.86 100.00 66.00 75.71 74.29 74.29 62.50 0.00 37.65 60.00 66.00 73.50 74.26 69.45 56.82 0.00 37.54 21.07
Independent Phind-CodeLlama-34B-v2 74.00 87.14 68.57 71.43 64.00 68.57 61.43 48.57 12.50 0.00 25.88 11.43 63.89 67.38 61.30 46.96 12.50 0.00 25.88 2.71
Independent gpt-3.5-turbo-1106 90.00 88.57 92.86 71.43 78.00 81.43 72.86 42.86 12.50 0.00 64.71 0.00 77.80 80.80 72.86 42.64 12.50 0.00 64.71 0.00
Average (Independent) 79.70 62.89 21.78 61.76 17.43
Incremental WizardCoder-15B-V1.0 70.00 74.29 48.57 28.57 46.00 51.43 35.71 17.14 0.00 0.00 0.00 0.00 45.31 50.84 35.68 16.62 0.00 0.00 0.00 0.00
Incremental deepseek-coder-6.7b-instruct 72.00 90.00 98.57 91.43 68.00 72.86 82.86 68.57 0.00 0.00 70.59 34.29 67.51 71.89 82.15 64.30 0.00 0.00 70.59 8.12
Incremental deepseek-coder-33b-instruct 68.00 77.14 92.86 85.71 54.00 74.29 71.43 54.29 55.00 0.00 14.12 0.00 54.00 73.99 71.38 49.81 49.08 0.00 14.02 0.00
Incremental Phind-CodeLlama-34B-v2 72.00 87.14 67.14 54.29 62.00 72.86 60.00 42.86 12.50 0.00 22.35 5.71 61.99 72.05 59.14 41.01 12.50 0.00 21.87 2.53
Incremental gpt-3.5-turbo-1106 72.00 85.71 90.00 91.43 62.00 80.00 71.43 68.57 10.00 0.00 0.00 25.71 61.50 79.80 71.43 64.84 10.00 0.00 0.00 8.10
Average (Incremental) 75.84 60.81 12.51 59.76 9.84
Average 82.42 65.34 23.08 64.15 18.22
@@
Generally, the average Completion@1 of all LLMs on JavaBench was 82.42%, with variances when different synthesis strategies were applied. Considering a finer granularity (class-wise), the average Compilation@1 and Pass@1 were around 65% (65.34% and 64.15%, respectively). The best scores were achieved using the holistic strategy with 91.73% Completion@1, 72.33% Compilation@1, and 70.92% Pass@1. Among all LLMs, DeepSeek-Coder-33b performed the best, followed by gpt-3.5-turbo and DeepSeek-Coder-6.7b.
@@
[style=MyFrame] Finding 1: The best average of 91.73%, 72.33%, and 70.92% Completion@1, Compilation@1, and Pass@1, respectively, could be achieved on JavaBench over the studied LLMs. The Top-3 performing LLMs were DeepSeek-Coder-33b, gpt-3.5-turbo, and DeepSeek-Coder-6.7b among the studied LLMs.
Comparing the synthesis strategies, holistic synthesis was generally better than independent and incremental among all LLMs, and the declines could be significant in some cases. For example, WizardCoder dropped 51.43% (80.00% - 28.57%) Completion@1, dropped 20.00% (37.14% - 17.14%) Compilation@1 and 19.52% (36.14% - 16.62%) Pass@1 on P4. Other LLMs also observed similar drops when switching from holistic to independent or incremental synthesis. Although there were occasional cases where incremental or independent synthesis brought improvements, those were sporadic events, and the improvement was subtle, e.g., Completion@1 of WizardCoder improved 4.29% (74.29% - 70.00%) on P2. We also noticed that this observation was different from the observation of the previous work (Du et al., 2023) where they observed open-sourced LLMs performed better using independent synthesis than holistic synthesis. This could be because we used different programming languages (Python VS. Java) and code granularities (Class-level VS. Project-level).
@@
Necessity of Finer-grained Evaluation Granularity. LLMs performed similarly on four projects (P1-P4), with P1 and P3 slightly better than P2 and P4. In particular, the class-wise scores were always better than test-wise scores, with a gap up to 49.92% (= 59.76%-9.84%). This was in line with our claim in Section 3.2.1: finer-grained granularity can capture more nuanced successful cases. We also calculated the Pass@1 under project-wise granularity and found that none of the projects can be correctly completed, yielding all-0 results under all settings.
@@
In addition, we increased the size of $k$ from 1 to 5 to investigate the improvements brought by more trials. The detailed experiment results are omitted due to space limits and can be found on our website (Section 8). Overall, the average Completion/Compilation/Pass@5 under holistic strategy across all LLMs are 97.21%(+5.48%), 84.43%(+12.10%), 84.43%(+13.51%) at the class-wise granularity; 94.21%(+5.48%), 51.23%(+16.28%), 48.24%(+20.84%) at the test-wise granularity. The Pass@5 in project-wise granularity is still all-0s under all settings. {mdframed}[style=MyFrame] Finding 4: Increasing $k$ yields further improvement. The best average test-wise Pass@5 in JavaBench is 48.24%. Compared with 90.93% achieved by undergraduate students in project-wise evaluation (Section 2.2.3), LLMs’ capability in Java project-level programming still has much room to improve.
@@ 4.2. RQ2: Impact of Context Selection
Selected WizardCoder-15B-V1.0 98.00 70.00 94.29 80.00 64.00 58.57 71.43 37.14 17.50 0.00 31.76 5.71 63.99 56.87 70.19 36.14 16.51 0.00 31.40 5.03
Selected deepseek-coder-6.7b-instruct 82.00 90.00 100.00 100.00 64.00 71.43 82.86 71.43 15.00 0.00 70.59 54.29 62.67 70.83 82.55 62.21 14.29 0.00 70.59 10.55
Selected deepseek-coder-33b-instruct 100.00 95.71 100.00 100.00 68.00 81.43 77.14 85.71 60.00 18.18 29.41 71.43 68.00 80.49 76.78 82.04 60.00 8.74 29.41 34.34
Selected Phind-CodeLlama-34B-v2 96.00 88.57 91.43 80.00 86.00 68.57 80.00 57.14 80.00 0.00 62.35 20.00 85.81 67.54 79.66 54.58 58.06 0.00 62.32 8.42
Selected gpt-3.5-turbo-1106 90.00 84.29 100.00 94.29 86.00 77.14 84.29 74.29 75.00 0.00 70.59 17.14 85.87 75.98 83.90 72.24 62.29 0.00 70.59 5.40
Average (Selected) 91.73 72.33 34.95 70.92 27.40
Maximum WizardCoder-15B-V1.0 98.00 70.00 88.57 54.29 64.00 58.57 58.57 14.29 37.50 0.00 21.18 0.00 63.97 57.66 57.85 13.29 34.12 0.00 21.18 0.00
Maximum deepseek-coder-6.7b-instruct 94.00 91.43 100.00 88.57 80.00 70.00 71.43 80.00 75.00 0.00 29.41 68.57 80.00 68.38 71.24 68.06 69.12 0.00 29.41 13.39
Maximum deepseek-coder-33b-instruct 100.00 95.71 100.00 100.00 74.00 80.00 75.71 68.57 67.50 18.18 45.88 25.71 74.00 79.36 75.63 62.23 67.50 13.64 45.67 10.72
Maximum Phind-CodeLlama-34B-v2 100.00 92.86 97.14 88.57 76.00 62.86 74.29 54.29 70.00 0.00 58.82 20.00 76.00 61.62 74.09 50.62 64.02 0.00 58.82 7.69
Maximum gpt-3.5-turbo-1106 94.00 85.71 100.00 94.29 68.00 68.57 64.29 62.86 55.00 0.00 36.47 22.86 67.56 66.07 63.88 59.77 51.47 0.00 36.47 7.78
Average (Maximum) 91.66 66.31 32.60 64.56 26.55
Minimum WizardCoder-15B-V1.0 100.00 78.57 90.00 77.14 18.00 30.00 48.57 0.00 0.00 0.00 5.88 0.00 18.00 29.22 48.52 0.00 0.00 0.00 5.88 0.00
Minimum deepseek-coder-6.7b-instruct 80.00 92.86 100.00 85.71 56.00 15.71 57.14 22.86 0.00 0.00 0.00 0.00 55.89 15.69 57.14 19.14 0.00 0.00 0.00 0.00
Minimum deepseek-coder-33b-instruct 100.00 91.43 98.57 82.86 48.00 22.86 65.71 11.43 12.50 0.00 37.65 0.00 48.00 22.58 64.98 5.21 12.50 0.00 37.43 0.00
Minimum Phind-CodeLlama-34B-v2 96.00 90.00 97.14 88.57 66.00 47.14 57.14 8.57 7.50 0.00 0.00 0.00 65.69 46.33 57.14 8.02 7.50 0.00 0.00 0.00
Minimum gpt-3.5-turbo-1106 98.00 85.71 95.71 91.43 66.00 40.00 62.86 20.00 12.50 0.00 0.00 0.00 65.94 39.85 62.86 19.27 12.50 0.00 0.00 0.00
Average (Minimum) 90.99 38.20 3.80 37.47 3.79
Overall Average 91.46 58.95 23.78 57.65 19.25
@@
From Table 6, it is clear that among three context settings, the selected context yield the best overall results, with 70.92% (class-wise) and 27.40% (test-wise) Pass@1, which echos the results in Table 5. Though maximum and minimum context achieve similar Completion@1 (i.e., 91.66% and 90.99% compared with 91.73%), the performance of using these two contexts on the subsequent metrics (i.e., Compilation@1 and Pass@1) is not as good as the selected context.
@@
[style=MyFrame] Finding 5: The selected context is the best setting on JavaBench, resulting in 70.92% (class-wise) Pass@1.
@@
By combining Table 6 and Figure 4, we can make two observations. First, more context is not always a benefit. For example, a dramatic 54.83% (= 69.12%-14.29%) increase in test-wise Pass@1 is achieved by DeepSeek-6.7b when switching the context from selected to maximum context in P1, while a 23.62% (= 34.34%-10.72%) drop of test-wise Pass@1 in P4 can be observed in DeepSeek-33b. Second, only providing the class to be completed is insufficient without dependencies. We can see from Table 6 that the Pass@1 in the test-wise granularity is almost all zeros under the minimum context setting, meaning that it is almost impossible to generate project-level code that can pass test cases. In addition, the selected context includes only the method signatures (as explained in Section 3.1.1), which turned out to be more effective than the maximum context setting.